OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 7 and the last term is -105?

OpenStudy (callisto):

Sum of AS= (first term +last term) x (no of terms) /2 = [7+ (-105)] x 17 /2 Can you solve it?

OpenStudy (anonymous):

-833..

OpenStudy (callisto):

I got the same answer as you.

OpenStudy (anonymous):

THIS VERY COMPLETE . . . \[n=17,a=7,U _{n}=-105\] So \[S _{n}=(n/2)(a+U _{b})=(17/2)(7-105)=17*(-98)=-833\]

OpenStudy (anonymous):

I mean \[S _{n}=(n/2)(a+U _{n})\]